OpenStudy (anonymous):

A club near you is offering judo classes which meet once a week. The classes cost $105 for 15 weeks or $170 for 26 weeks. How much can you save per class by taking the 26 week classes instead of the 15 weeks? Less than $1 Less than $5 Less than $10 <--This is what I got More than $10 Please help

OpenStudy (anonymous):

105 / 15 =? 170/ 26 =?

OpenStudy (anonymous):

\[170\div26\]

OpenStudy (anonymous):

\[105\div15\]

OpenStudy (anonymous):

now, subtract the 2nd one from the first, and that's your answer

OpenStudy (anonymous):

$105 / 15 = $7 $170/26 = $6.54 so the differnce of 7 dollars to 6.54 dollars is less than 1 dollar differnce

OpenStudy (anonymous):

That's what I got too.

OpenStudy (anonymous):

sorry, i thought in your post you got "less than $10"